Between last year and this year, the CPl In Blueland rose from 100 to 110 and the CPI In Redland rose from 100 to 105. Blueland's currency unit, the blue, was worth $1.00 (U.S.) last year and is worth $O.90 (U.S.) this year. Rediand's currency unit, the red, was worth S0.50 (U.S.) last year and Is worth $0.45 (U.S.) this year. Consider Blueland as the home country a. Calculate Blueland's nominal exchange rate with Redland Instructions: Enter your response rounded to one decimal place. Last year. red/blue This year:red/blue The percentage change In Blueland's nominal exchange rate from last year to this year Is: Instructions: Enter your response as a whole number. Be certaln to enter "O" If required. b. Calculate Blueland's real exchange rate with Redland Instructions: Enter your response rounded to two decimal places Last year red/blue This yearred/blue
The percentage change In Blueland's nominal exchange rate from last year to this year is: Instructions: Enter your response as a whole number. Be certaln to enter "O If required b. Calculate Blueland's real exchange rate with Redland. Instructions: Enter your response rounded to two decimal places. Last yearred/blue This year: The percentage change In Blueland's real exchange rate with Redland from last year to this year ls: Instructions: Enter your response rounded to two decimal places. Be certaln to enter "O" If required. red/blue C. Relative to Rediand, you expect Blueland's exports to be [Click to select by these changes in exchange rates.
Answer #1

a. Nominal Exchange Rate of Blueland with Redland

This Year = Blue/Red = 0.90/0.45= 2 or Blue = 2*Red

Last Year = Blue/Red = 1/0.50= 2 or Blue = 2*Red

The Percentage Change is zero as there is no change in Nominal Exchange from last year to this year in Blueland

b. Real Exchange Rate = Nominal Exchange Rate *(Pf / Pd)

This Year = 2*(105/110) = 1.909

Last Year = 2*(100/100)= 2

The Percentage Change is given below

(1.909 - 2)*100/2 = 4.55%